OpenStudy (erinweeks):

Someone please help!! MEDAL?(:

OpenStudy (erinweeks):

In an experiment, a petri dish with a colony of bacteria is exposed to cold temperatures and then warmed again. Time (hours) 0 1 2 3 4 5 6 Population (1000s) 5.1 3.03 1.72 1.17 1.38 2.35 4.08

OpenStudy (erinweeks):

a. Find a quadratic model for the data in the table.

OpenStudy (erinweeks):

b. Use the model to estimate the population of bacteria at 9 hours.

OpenStudy (anonymous):

@zepdrix

Directrix (directrix):

Look at the attachment below to view how the data was entered. Check out the quadratic regression online calculator at: http://www.had2know.com/academics/quadratic-regression-calculator.html @ErinWeeks

OpenStudy (nikato):

@ErinWeeks u can set up 3 system of equations to solve for ur A, B, and C in ur quadratic equation becuz a quadtratic equation is in the form Ax^2+Bx+c=y so...u can have \[5.1=a(0^{2})+b(0)+c\]

OpenStudy (nikato):

\[1.72=a(2^{2})+b(2)+c\]

OpenStudy (nikato):

\[1.17=a(3^{2})+b(3)+c\]

OpenStudy (nikato):

and all u ahve to do is solve for a and b

OpenStudy (erinweeks):

now why am I just solving for those specific calculations?

OpenStudy (nikato):

actually, u dont have to, i just picked those. u can pick any other 3 if u want to

OpenStudy (erinweeks):

It doesnt matter which ones I have to solve for or is it like I can choose any of those calculations and for one of them I just have to solve for A & B?

OpenStudy (nikato):

no, u need all three of those to solve. its call a system of equations. this is what u do to solve for more than 1 varibale. when i said choose any 3. i meant any pairs of ur x and y from ur data

OpenStudy (nikato):

u have ur basic quadratic equation right? \[a ^{2}+bx+c=y\]

OpenStudy (erinweeks):

ok we can use those three if itll be enough to get my question right... what do we start off with

OpenStudy (nikato):

with the first equation, after simplifying, u should see that c=5.1. ok?

OpenStudy (erinweeks):

ok how did I get it to C=5.1

OpenStudy (nikato):

see, my first equation. this si what u get after simplifying